OpenStudy (anonymous):

3p - 1 = 5(p - 1) - 2(7 - 2p)

jigglypuff314 (jigglypuff314):

3p - 1 = 5(p - 1) - 2(7 - 2p) distribute 3p - 1 = 5p - 5 - 14 + 4p add like terms 3p - 1 = 9p - 19 subtract 3p from both sides -3p -3p -1 = 12p - 19 add 19 to both sides +19 +19 18 = 12p divide both sides by 12 p = 18/12 = (simplify)

OpenStudy (anonymous):

much appreciated

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

3p-1=5p-5-14-4p 3p-1=5p-9-4p 3p+8=5p-4p 3p+8=1p 4=p

jigglypuff314 (jigglypuff314):

you forgot to distribute the negative in the first step @sportcraze0

OpenStudy (anonymous):

Thanks. I looked at the problem wrong.